N is a five- digit positive integer. A six-digit integer P is constructed by placing a 1 at the right hand end of N. A second six digit Q is constructed by placing a 1 at the left hand end of N. If P is three times Q, determine the value of N.
Found 2 solutions by Edwin McCravy, w_parminder:
Answer by Edwin McCravy(20059)

N is a five- digit positive integer. A six-digit integer P is constructed by placing a 1 at the right hand end of N. A second six digit Q is constructed by placing a 1 at the left hand end of N. If P is three times Q, determine the value of N.
      P = 10N + 1
      Q = 100000 + N

      P = 3Q

10N + 1 = 3(100000 + N)
10N + 1 = 300000 + 3N
     7N = 299999
      N = 42857

Edwin


Answer by w_parminder(53)


Given 5-digit number = N
A.T.Q.
P = 10 x N + 1
= 10N + 1
Q = 100000 + N
Also P = 3 x Q
So 10N + 1 = 3(100000 + N)
10N + 1 = 300000 + 3N
10N - 3N = 300000 - 1
7N = 299999
N = 299999/7
N = 42857